OpenStudy (anonymous):

Find the point on the terminal side of θ = negative three pi divided by four that has an x coordinate of -1.

OpenStudy (e.mccormick):

\(\theta=-\dfrac{3\pi}{4}\) Did you sketch that? It could help here.

OpenStudy (e.mccormick):



OpenStudy (anonymous):

Thank you. I'm just stuck

OpenStudy (e.mccormick):

Do you know what the angle \(\dfrac{\pi}{4}\) is in degrees? It will help solve this.

OpenStudy (e.mccormick):